OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). cos x = sin x

OpenStudy (anonymous):

You can do this algebraically/geometrically. Algebraically: Start at x=0 is cos(0)=sin(0) ? No, cos(0)=1 and sin(0)=0 move on to other values within [0, 2pi) Geometrically graph them and see where the graphs intersect.

OpenStudy (anonymous):

Here is an easy one: When x=45 degrees we have cos(x)=sin(x) because cos(45)=sin(45)=.707106...

OpenStudy (anonymous):

Or you can look at the unit circle and look at which angles the x coordinate (cosine) equals the y coordinate (sine).

OpenStudy (anonymous):

When you look at the interval [0, 2pi) you are looking at all the angles from the x-axis all the way around until you end up in the x-axis again (when you make the 360). You should notice that at 45degrees (or pi/4) and at 225 degrees (or 5pi/4) cosx=sinx